OpenStudy (anonymous):

There were 44,000 people at a ball game in Los Angeles. The day's receipts were $ 369,000. How many people paid $ 13.00 for reserved seats and how many paid $ 6.00 for general admission?

OpenStudy (anonymous):

People who paid 13 = p paid 6 = c p + c = total of people = 44,000 p * 13 + c * 6 = total receipt = 369,000 p + c = 44,000 - c = 44,000 - p 13 p + 6 c = 369,000 13 p + 264,000 - 6 p = 369,000 7p = 105,000 p = 15,000 c = 29,000